OpenStudy (anonymous):

Find the missing number in the following series: 4 ? 17 51 56 168

ganeshie8 (ganeshie8):

12 two things happening here 1. next number is 3 times the current number. 2. next number is 5 units more than the current number lets start from end : 168 = 56 x 3 56 = 51 +5 51 = 17 x 3 17 = 12 + 5 12 = 4x3